# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

### v5

@@ 8 Theoretical Context and Limitations
Limitations of the Model
While the model presented here offers a novel thermodynamic lens through which to understand the emergence of natural selection, several limitations must be acknowledged. These do not detract from the value of the conceptual framework but highlight areas for refinement, expansion, and empirical exploration.
Simplification of Chemical Complexity
This model abstracts chemical reactions into energetically discrete events, using parameters such as $x_{i}$, $r_{i}$, and $s_{i}$ to represent energy thresholds, releases, and storage. In reality, reactions involve complex molecular dynamics, intermediate states, and environmental feedbacks. No specific chemistry is modeled, and thus, while the logic is general, its direct applicability to known prebiotic chemistries remains to be validated.
Environmental Energy Model is Limited
The energy input function $z(t)$ is treated abstractly and assumed to vary arbitrarily over time. While this reflects prebiotic environments (e.g., thermal cycling, day-night radiation), no specific environmental model is encoded. Incorporating realistic environmental data or simulating z(t) from known planetary conditions could help test the plausibility of different prebiotic scenarios.
Related Works and Theories
Autocatalytic Sets and Organizational Closure
@@
Stuart Kauffman’s theory of autocatalytic sets (Kauffman, 1993) laid the foundation for viewing life as a network of reactions that sustain and propagate themselves. His work framed emergence not as a consequence of genetic control but as a product of combinatorial chemistry and mutual reinforcement. Our model complements this by suggesting that not all autocatalytic chains persist equally, only those aligned with dynamic energy availability survive.
Dissipative Structures and Far-From-Equilibrium Systems
@@
Ilya Prigogine’s concept of dissipative structures (Prigogine and Nicolis, 1977) posits that ordered systems can arise spontaneously in non-equilibrium environments through energy dissipation. Our framework draws from this principle by proposing that prebiotic systems capable of balancing energetic inflow and reaction requirements naturally self-stabilize, forming precursors to selection mechanisms.
Wet–dry cycles cause nucleic acid monomers to polymerize into long chains
@@
Song et al.’s work on nucleotide polymerization (Song et al, 2019) investigates how nucleic acid monomers, when subjected to repeated wet–dry cycles under conditions simulating volcanic hot springs, can spontaneously polymerize into long chains. This study demonstrates that RNA and DNA–like oligomers can form in the absence of enzymes, highlighting a plausible non–enzymatic pathway for early genetic polymers. Their findings support the idea that environmental cycling, particularly dehydration–rehydration dynamics, can drive condensation reactions that would otherwise be thermodynamically unfavorable in continuously aqueous conditions.
Generation of oligonucleotides under hydrothermal conditions by non-enzymatic polymerization
@@
Deguzman et al’s work on hydropthermal conditions (DeGuzman et al, 2019) studies prebiotic synthesis under fluctuating hydration that support the idea of nonequilibrium thermodynamic cycling can create and stabilize chemical complexity, without the need for enzymes or genetic templates.
Dissipative Adaptation
@@
Jeremy England’s work on dissipative adaptation (England, 2013) formalized how driven systems statistically evolve to dissipate energy more efficiently. We align with this view, proposing that selection arises not from genetic memory, but from energy-aligned structural persistence, chains that survive fluctuations in $z(t)$ naturally dominate the prebiotic landscape.
Dynamic Kinetic Stability
@@
Addy Pross’s theory of dynamic kinetic stability (Pross, 2012) reframes life’s origin as a transition from thermodynamic to kinetic dominance, that is, from stable molecules to stable replicative processes. Our model precedes this transition, exploring the energetic logic by which certain reaction chains become stable enough to eventually support replication.
The Hydrophobic Effect and Amphiphilic Self-Assembly
@@
The hydrophobic effect explains the spontaneous organization of amphiphilic molecules in aqueous environments. First formalized in detail by Charles Tanford (Tanford, 1980), the effect arises from water’s tendency to exclude non-polar groups in order to preserve its hydrogen-bonding network. When hydrophobic molecules or molecular segments are introduced into water, they disrupt the local structure of the solvent, resulting in an entrop…
This behavior is central to many biological structures, including lipid membranes, and forms the physical basis for compartmentalization in the absence of genetic instruction.
Novelty of the Present Work
While these foundational theories address persistence, emergence, and order, they rarely formalize persistence in terms of a reaction viability inequality under dynamically shifting energy constraints. This work contributes a simplified but generalizable framework in which selection arises not from heredity, but from energy dynamics and reaction compatibility. It provides a thermodynamically grounded filter, one that operates even before life, as typically defined, begins.
9 Model Flexibility and Testing

### v7

@@ 7 Theoretical Context and Limitations
7.1 Limitations of the Model
7.1.1 Simplified Chemical Representation
The model treats reactions as discrete energetic steps parameterized by $x_{i}$, $r_{i}$, and $s_{i}$. This abstraction omits mechanistic detail—such as intermediates, catalysis, and solvent-structure couplings—in favor of generality. Although this enables broad applicability, its mapping to specific prebiotic chemistries must be established empirically. Experimental and computational studies will be required to determine realistic parameter ranges for early Earth or analogous environments.
7.1.2 Environmental Energy Abstraction
The environmental input term $z(t)$ is modeled generically to encompass diverse prebiotic energy sources (thermal cycling, photonic flux, geothermal gradients). While appropriate for a general framework, it does not encode any specific planetary setting. Future work could incorporate empirical energy profiles—diurnal intervals, geothermal temperature curves, or seasonal variations—either analytically or through simulation. Extending $z(t)$, $x_{i}$, and $r_{i}$ to depend explicitly on temperature or other environmental variables (e.g., Arrhenius-like behavior) would allow more detailed comparisons without altering the framework’s core structure.
7.1.3 Stored Energy and Buffering
The model treats stored energy as a structural or chemical reserve capable of buffering fluctuations in $z(t)$. This captures the intuitive idea that systems with internal energy stores can survive transient loss events. However, the biophysical mechanisms that might instantiate S(t)S(t)—such as polymerized energy carriers or metastable structures—remain to be explored in real chemical systems. Quantifying the magnitude and decay of such stores is an important next step for connecting the inequality to experiment.
7.1.4 Entropy and Spatial Extensions
The augmented inequality incorporates entropic and spatial terms ($\Delta S_{local}$, $\Delta S_{env}$, $D(x,t)$), enabling treatment of self-assembly, diffusion-limited organization, and compartmentalization. These extensions broaden the model’s physical realism but also introduce additional parameters that must be experimentally constrained. The current formulation treats these contributions phenomenologically rather than mechanistically, and follow-up work will be required to map these abstract terms onto measurable quantities in specific chemical systems.
7.1.5 Reaction Networks and Topology
The present implementation considers linear reaction chains, whereas real prebiotic systems likely involved branching pathways, redundancy, and feedback. Extending the framework to network topology would clarify how resilience emerges from alternative pathways and whether autocatalytic or DKS-like organizations appear naturally within the persistence criterion. This represents a natural avenue for future generalization.
7.1.6 Thermodynamic Origins of Selection
The model posits that selection precedes heredity: structures that maintain viability under fluctuating energy inputs persist, while others collapse. This thermodynamic filtering aligns with known features of dissipative systems, autocatalytic organization, and energy-entropy-compatible self-assembly. The resulting picture is probabilistic rather than discrete: abiogenesis emerges as a gradual transition in which open systems explore chemical space while energy flow retains those configurations most compatible with the local environment.
These persistence inequalities therefore offer a pre-replicative substrate on which later Darwinian constraints—such as those formalized by Kolchinsky—might act once replication appears. Verifying the predictions made here would establish whether energy-structure coherence provides the physical preconditions for natural selection, bridging the gap between purely thermodynamic organization and the efficiency-driven competition characteristic of life.
7.2 Related Works and Theories
7.2.1 Autocatalytic Sets and Organizational Closure
@@
Stuart Kauffman’s theory of autocatalytic sets [1] laid the foundation for viewing life as a network of reactions that sustain and propagate themselves. His work framed emergence not as a consequence of genetic control but as a product of combinatorial chemistry and mutual reinforcement. Our model complements this by suggesting that not all autocatalytic chains persist equally, only those aligned with dynamic energy availability survive.
7.2.2 Dissipative Structures and Far-From-Equilibrium Systems
@@
Ilya Prigogine’s concept of dissipative structures [12] posits that ordered systems can arise spontaneously in non-equilibrium environments through energy dissipation. Our framework draws from this principle by proposing that prebiotic systems capable of balancing energetic inflow and reaction requirements naturally self-stabilize, forming precursors to selection mechanisms.
7.2.3 Wet–dry cycling and non-enzymatic polymerization
@@
Song et al.’s work on nucleotide polymerization [9] investigates how nucleic acid monomers, when subjected to repeated wet-dry cycles under conditions simulating volcanic hot springs, can spontaneously polymerize into long chains. This study demonstrates that RNA and DNA-like oligomers can form in the absence of enzymes, highlighting a plausible non-enzymatic pathway for early genetic polymers. Their findings support the idea that environmental cycling, particularly dehydration-rehydration dynamics, can drive condensation reactions that would otherwise be thermodynamically unfavorable in continuously aqueous conditions.
7.2.4 Generation of oligonucleotides under hydrothermal conditions by non-enzymatic polymerization
@@
Deguzman et al’s work on hydrothermal conditions [22] studies prebiotic synthesis under fluctuating hydration that support the idea of nonequilibrium thermodynamic cycling can create and stabilize chemical complexity, without the need for enzymes or genetic templates.
7.2.5 Dissipative Adaptation
@@
Jeremy England’s work on dissipative adaptation [2] formalized how driven systems statistically evolve to dissipate energy more efficiently. We align with this view, proposing that selection arises not from genetic memory, but from energy-aligned structural persistence, chains that survive fluctuations in $z(t)$ naturally dominate the prebiotic landscape.
7.2.6 Dynamic Kinetic Stability
@@
Addy Pross’s theory of dynamic kinetic stability [3] reframes life’s origin as a transition from thermodynamic to kinetic dominance, that is, from stable molecules to stable replicative processes. Our model precedes this transition, exploring the energetic logic by which certain reaction chains become stable enough to eventually support replication.
7.2.7 The Hydrophobic Effect and Amphiphilic Self-Assembly
@@
The hydrophobic effect explains the spontaneous organization of amphiphilic molecules in aqueous environments. First formalized in detail by Charles Tanford [4], the effect arises from water’s tendency to exclude non-polar groups in order to preserve its hydrogen-bonding network. When hydrophobic molecules or molecular segments are introduced into water, they disrupt the local structure of the solvent, resulting in an entropic penalty. To minimize this disruption, water reorganizes itself into a more ordered shell around the solute. However, when multiple hydrophobic groups cluster together, the overall surface area exposed to water is reduced, thereby releasing structured water molecules back into the bulk and increasing the total entropy of the system.
This process underlies the self-assembly of amphiphilic molecules-those with both hydrophilic and hydrophobic domains, into micelles, bilayers, and vesicles. The hydrophobic tails aggregate to minimize exposure to water, while the hydrophilic heads remain solvent-accessible. As a result, structurally coherent assemblies emerge without any external guiding information.
Hydrophobic Theories of Self-Organization
A number of theoretical models have been developed to explain the behavior of amphiphilic molecules and the emergence of biologically significant structures such as micelles, bilayers, and folded proteins, based on hydrophobic interactions.
Micelle formation
Micelle formation represents a foundational example of hydrophobic self-assembly. Amphiphilic molecules such as fatty acids or detergents spontaneously aggregate in aqueous environments to form micelles-spherical structures in which the hydrophobic tails are shielded from water by an outer layer of hydrophilic headgroups. This process minimizes the entropic cost of disrupting water’s hydrogen-bonding network.
Thermodynamic models, such as those developed by Maibaum et al (2004), have demonstrated how micelle formation can be simulated using coarse-grained particles interacting via effective hydrophobic potentials [23]. Their work provides a statistical mechanical basis for the emergence and reorganization of micelles under different thermodynamic constraints. However, these models generally assume that the amphiphilic molecules are already present in optimal concentrations and that micelle formation is driven purely by equilibrium considerations.
Lipid Bilayers and Bilayer Formation
Brannigan et al (2005). introduced an implicit-solvent model in which coarse-grained lipid-like amphiphiles self-assemble into bilayer structures [24]. Their work captures membrane fluidity and mechanical flexibility through effective hydrophobic interactions, providing a simplified physical basis for bilayer formation. However, their model begins with designed molecules and assumes favorable interactions, rather than offering an explanation for why such structures would emerge prebiotically. Hence the thermodynamic persistence model can be viewed as a complementary framework, offering criterion by which certain amphiphilic molecules, among many possible prebiotic variants, would be preferentially retained under fluctuating energetic conditions.
Interface-Sensitive Assembly Behavior
Patel et al (2011). extended the Lum-Chandler-Weeks theory to show how hydrophobic effects at extended surfaces can catalyze assembly and influence water structure [25]. This framework contributes a deeper understanding of how surface geometry and water fluctuations modulate hydrophobic assembly. These environmental variables could be integrated into the viability model to evaluate how structured surfaces affect reaction persistence, pointing to a possible bridge with interface-sensitive assembly behavior.
The Energy Funnel Model of Protein Folding
Hummer et al (1998) developed a statistical mechanical theory for hydrophobic solvation at the molecular scale [26], revealing how entropy and solvent structuring contribute to protein folding. This was built on to propose the energy funnel model of protein folding [27], resolving the Levinthal paradox by showing how hydrophobic collapse drives efficient exploration of conformational space. The persistence inequality model does not directly reference model folding, but it may explain why certain polymers with folding-prone sequences might persist longer than others, thus serving as a precondition for the emergence of foldable biomolecules.
7.3 Novelty of the Present Work
While these foundational theories address persistence, emergence, and order, they rarely formalize persistence in terms of a reaction viability inequality under dynamically shifting energy constraints. Existing frameworks – from dissipative structure theory [7] to maximum entropy production principles [28] and dynamic kinetic stability [3] - successfully describe how organized states can arise and be maintained in driven environments. However, these approaches typically operate at macroscopic or statistical levels and do not provide a local, quantitative criterion for determining whether a specific prebiotic reaction system will continue, collapse, or transition under fluctuating external conditions. The present work introduces such a criterion in the form of a reaction viability inequality that links environmental energy input, internal energetic costs, and entropy-related penalties to the time-dependent persistence of a reaction network.
This inequality-based formulation is distinct in three respects. First, it identifies selection-like behaviour prior to heredity or replication by showing how differential persistence emerges purely from thermodynamic compatibility with the driving environment. This contrasts with frameworks such as England’s dissipative adaptation, which presuppose replication-like transitions [2]. Second, it provides a minimal but generalizable filter that applies at the level of individual reactions, small networks, or mesoscale assemblies, complementing global dissipation-optimization views that operate at much larger scales, such as biospheric entropy production models [29]. Third, it translates the thermodynamic constraints on persistence into experimentally testable conditions, enabling direct evaluation in systems such as wet–dry polymerization cycles and lipid-based protocells where environmental driving determines structural longevity [30].
In this way, the framework does not attempt to replace existing non-equilibrium theories, but instead provides a local, mechanistic bridge between environmental driving and the differential survival of reaction networks – a domain where current formalisms offer descriptive insight but no operational criterion for viability.
8 Implications, Predictions, and Falsifiability
8.1 Theoretical Outlook
The framework presented, provides a means of translating this conceptual model into a probabilistic structure applicable across diverse environmental conditions. The reaction viability inequality, extended via $y^{\prime}(t)$, allows for the estimation of persistence under spatial and energetic modulation. This yields not only a theoretical basis for selection-like behavior in prebiotic chemistry but also a potential comparative lens across planetary systems.
Continued refinement of this approach, through simulation of reaction networks, laboratory analogs of compartmentalized cycling systems, and the thermodynamic profiling of planetary bodies, may illuminate the constraints and pathways most conducive to the emergence of persistent reaction architectures. In this light, the framework serves not only as a theoretical advance but as a scaffold for the systematic comparison of environments where the thermodynamic roots of natural selection may have taken hold.
Moreover, pairing the model with Urability, the environmental pre-screening framework [21], which assesses whether a planetary setting possesses the thermodynamic and chemical conditions necessary for prebiotic complexity, provides a two-tiered approach by first evaluating urability, then applying the Thermodynamic Abiogenesis Likelihood Model. This provides a robust, scalable method for identifying environments most conducive to abiogenesis.
The expanded test case scenarios now incorporated into this work reinforce the models generality. They demonstrate that the persistence inequality can be applied not only to abstract reaction networks, but to real physical systems, such as folding oligomers, membrane dynamics, and compartmental reactions. These testable extensions highlight the Thermodynamic Abiogenesis Likelihood Model predictive versatility and provide a concrete roadmap for future validation. Crucially, they show that persistence inequality is not a metaphorical selection mechanism, but a quantitative framework grounded in measurable energy flows.
8.2 Predictions and Falsifiability
The experimental and simulation scenarios described in Sections 8-10 collectively serve a single purpose: to determine whether the reaction viability inequality and its augmented persistence function accurately describe how chemical systems maintain structure under fluctuating energy input. Each test therefore represents an opportunity for the model to be proven wrong.
8.3 Purpose of the suite
Together, the tests probe three linked claims of the thermodynamic persistence inequality model:
  1. 1.
    Persistence is governed by energetic alignment, not replication;
  2. 2.
    Entropy export and spatial confinement are the decisive modulators of structural stability; and
  3. 3.
    These same principles scale from molecular assemblies to planetary environments.
8.4 Test Scenarios
8.4.1 Entropy-Driven Self-Assembly in Amphiphilic Systems
A tractable and highly illustrative system for testing the persistence inequality model involves the self-assembly behavior of amphiphilic molecules in water, where structure formation is driven primarily by the hydrophobic effect. This effect arises because water molecules form hydrogen-bonded networks. When non-polar (hydrophobic) molecules are introduced, they disrupt these networks, lowering the system’s entropy. Water responds by ordering tightly around the intruding molecules. However, when many hydrophobic molecules are present, they tend to cluster together, freeing the surrounding water to return to a disordered state, thus increasing entropy.
To explore this within the persistence model framework, consider a series of aqueous systems each containing 18 mL of water (1 mole) and 0.01 mole of one of three solutes: sodium acetate, sodium octanoate, and sodium dodecanoate. These molecules differ only in the length of their hydrophobic tails. When air is gently introduced, observable differences arise. Sodium acetate, which lacks a hydrophobic tail, dissolves completely with no visible change. Sodium octanoate, with a moderate tail length, forms a transient bubble. Sodium dodecanoate, with a longer hydrophobic chain, forms a stable bubble and, under microscopic inspection, many vesicles.
These behaviors are predicted as a function of chain length and the system’s ability to manage entropy and energy flow. The longer tails drive stronger self-association, lowering diffusion ($D(x,t)$) and increasing structural resilience ($R_{n}$), while also allowing water to release more entropy as structured cages collapse. Thus, the entropy cost of forming vesicles is more than offset by entropy export to the environment:
$\Delta S_{net}=\Delta S_{env}-\Delta S_{local}>0$
leading to a reduced $\Phi(t)$ and positive $y^{\prime}(t)$.
We predict that structural persistence-bubbles, micelles, and vesicles emerge at a critical chain length $L_{c}$. Based on this system:
  • For $L<C8$, self-assembly is minimal; no structures persist.
  • At $L\approx C10$, vesicle formation becomes likely.
  • Between $C10$ and $C14$, vesicles are increasingly stable.
  • Beyond $C14$, solubility may decrease, limiting structure formation.
Tests
  • Context/Gap: Experimental studies show that fatty acid vesicles depend strongly on chain length for their stability, but no general thermodynamic criterion has been formulated for identifying the thresholds of persistence under fluctuating conditions.
  • Model Action: The thermodynamic persistence model translates amphiphile self-assembly into a balance between energy thresholds for aggregation and entropy penalties from solubility and dispersion.
  • Prediction (Expectation): Vesicles shorter than C8 will fail to persist, vesicles near C10 will form but remain fragile, vesicles between C10 and C14 will persist more reliably, and chains longer than C14 will become limited by poor solubility.
  • Falsification: If experiments show stable vesicles persisting well below C8 or well above C14 without confinement or catalysis, or if vesicle stability is largely independent of chain length, the framework would be falsified in this domain.
8.4.2 Surface-Catalyzed Reaction Viability
Mineral surfaces can influence the persistence of prebiotic reaction networks by modulating local energy dissipation, adsorption, and spatial retention. Within the persistence framework, the viability of a reaction sequence is evaluated using the condition
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)>0.$
Surface properties influence the loss term $\Phi(t)$ by altering water activity, reactant residence time, and the extent of molecular dispersal. Hydrophobic or structurally repetitive surfaces may restrict diffusion and lower environmental loss, whereas highly hydrated or reactive surfaces may increase $\Phi(t)$ through accelerated hydrolysis or rapid desorption.
To test this, we propose placing identical chemical mixtures on two surfaces: one smooth and water-repelling (hydrophobic), the other patterned and water-attracting (hydrophilic). Both are exposed to repeated wet-dry and heat intervals. If the prediction holds, the surface that supports a better energy balance, where $y^{\prime}(t)>0$, will see the reactions last longer.
This experiment would aim to demonstrate that even before enzymes or catalysts existed, the shapes and chemistry of early surfaces may have influenced chemical persistence, selecting not by function, but by physics.
The persistence inequality model predicts that structured surfaces will differentially support reaction persistence based on how they modulate energy dissipation:
  • On hydrophilic patterned surfaces: high water structuring increases $\Phi(t)$, reducing persistence for some reactions.
  • On hydrophobic flat surfaces: clustering may lower $\Phi(t)$, favoring reactions with minimal net energy waste.
  • Reaction networks aligned with local water organization will recur more reliably across energy intervals.
  • Surfaces that buffer thermal or hydration fluctuations will increase $y(t)$ enhancing persistence.
Tests
  • Context/Gap: Mineral and rock surfaces are widely invoked in prebiotic chemistry, yet existing models often treat them as passive scaffolds rather than active thermodynamic constraints on persistence.
  • Model Action: In the thermodynamic persistence model, surface effects are captured by the entropic-diffusive penalty term $\Phi(t)$, which reflects water structuring, clustering, and buffering of fluctuations.
  • Prediction (Expectation): Hydrophilic patterned surfaces should increase $\Phi(t)$ and reduce persistence, while hydrophobic surfaces should reduce $\Phi(t)$ and favor persistence. Surfaces that buffer hydration or thermal intervals should increase $y(t)$ and enhance persistence.
  • Falsification: If persistence is observed to be unaffected by surface hydrophobicity, structuring, or buffering capacity, with systems persisting equally across all surface types, then the framework would be falsified in this context.
8.4.3 Folding Bias in Random Oligomer Pools
Non-templated peptides and RNA oligomers produced in prebiotic settings would have exhibited wide variability in structural stability. Many sequences would remain extended and solvent-exposed, rendering them susceptible to hydrolysis, thermal degradation, or dispersal. Others would spontaneously adopt compact folds that reduce the effective surface area exposed to the environment, particularly by burying hydrophobic residues or bases within an interior core.
Within the thermodynamic persistence framework, structural compaction contributes directly to viability through its effect on the persistence inequality:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)>0$
Here, $y(t)$ represents energy stored or buffered, folding adds to this; $R_{n}$ is the net energy gained from any reaction (possibly zero here); and $\Phi(t)$ is energy lost to the environment through heat or degradation.
A folded molecule might survive better simply because it reduces $\Phi(t)$ and maintains $y(t)$, keeping the total $y^{\prime}(t)$ positive.
To test this idea, we can create a large batch of random RNA or peptide sequences in water. Then we expose them to simulated early-Earth conditions: heating and cooling, drying and re-wetting, maybe even light exposure. After several intervals, we check which sequences are still around.
The model predicts that many of the survivors will be sequences that naturally fold, those that form compact, protected shapes in water. Folding, in this case, becomes a kind of passive survival mechanism. This would suggest that even before biology had evolved to select useful sequences, nature was already filtering out unstable ones, simply based on their energy behavior in a rough environment.
Hence, foldable sequences will be statistically enriched among survivors in fluctuating environments:
  • Sequences that form hydrophobically collapsed folds exhibit higher $y(t)$ (energy retention).
  • Folding reduces exposure to destructive fluctuations, lowering $\Phi(t)$.
  • Non-folding sequences degrade or hydrolyze more quickly during thermal or hydration cycling.
  • Over time, the population shifts toward structurally compact, energy-buffering polymers.
Tests
  • Context/Gap: Prebiotic polymerization would have produced vast numbers of sequences, most of which are non-folding. Existing theories have not explained how foldable sequences could be selectively retained prior to heredity.
  • Model Action: Folding alters thermodynamic viability by raising $y(t)$ (through energy retention in compact states) and lowering $\Phi(t)$ (by reducing exposure to destructive fluctuations).
  • Prediction (Expectation): Foldable polymers should persist more reliably across intervals of hydration and heating, while non-foldable polymers should degrade or hydrolyze more rapidly. Over time, foldable sequences should be statistically enriched.
  • Falsification: If experiments on polymer libraries reveal no persistence differences between foldable and non-foldable sequences under repeated cycling, the framework would be invalidated in this domain.
8.4.4 Micelle-Embedded Reaction Selection
Amphiphilic molecules in aqueous solution can self associate into micelles from what is referred to as the hydrophobic effect, where water avoids oily parts of molecules. This process decreases the solvent exposed surface area and produces a protected hydrophobic core. Such structures can buffer environmental fluctuations by reducing the entropic and diffusive loss term $\Phi(t)$ and by stabilizing stored energy contributions to $y(t)$.
According to the persistence model, a micelle whose internal chemistry balances energy well, absorbing, storing, and releasing it in sync with the environment, should persist longer than one filled with wasteful or unstable reactions. The principle is summarized in the persistence inequality equation:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)$
Where $y(t)$ is the energy held inside the micelle (e.g., as chemical potential), $R_{n}$ is the net energy from internal reactions, and $\Phi(t)$ is energy dissipated (through heat or structural strain).
Micelles that keep $y^{\prime}(t)>0$, that is, they maintain more usable energy than they lose, should survive longer through intervals of drying, heating, or stress.
To test this, scientists could create two groups of micelles with the same membranes but different contents: one group with gentle, energy-buffered reactions inside, and one with inefficient, chaotic reactions. After exposing both to intervals of dehydration and rehydration, the persistence inequality model predicts that the first group will re-form more often and persist longer.
We predicts that micelles with internal reactions that support energy alignment will dominate over time:
  • Encapsulated reactions with steady, buffered energy flow increase $y(t)$ and reduce internal $\Phi(t)$.
  • Micelles containing inefficient or highly dissipative reactions rupture or fail to reform.
  • Micelle populations with aligned internal networks grow more numerous over intervals.
  • Selection acts on the reaction-micelle composite, not just the structure.
Tests
  • Context/Gap: Micelles are known to form spontaneously, but the impact of internal chemistry on their persistence under fluctuating conditions remains unclear.
  • Model Action: The thermodynamic persistence model incorporates encapsulated reactions into the viability inequality by coupling their energy dynamics to the host structure, modifying both $y(t)$ and $\Phi(t)$.
  • Prediction (Expectation): Micelles containing buffered internal reactions should maintain higher $y(t)$ and lower $\Phi(t)$, while micelles hosting dissipative reactions should rupture or fail to reform. Over time, selection will act on the reaction-micelle composite.
  • Falsification: If micelles persist equally regardless of whether they contain energy-aligned or dissipative reactions, the framework’s predictions would be contradicted.
8.4.5 Bilayer Reconfiguration Under Cycling Energy Input
Lipid bilayers can be modeled and simulated to understand how their structure depends on molecular properties like chain length and flexibility. The persistence model adds a new layer to this understanding: it suggests that bilayers which better regulate energy flow, absorbing and releasing it without collapsing, will persist longer under environmental cycles.
We express this idea by applying the simple inequality equation:
$y^{\prime}(t)=y(t)+R_{n}-\Phi(t)>0$
Here, $y(t)$ is energy retained in the bilayer system (e.g., mechanical stability, membrane potential); $R_{n}$ is energy added by any internal reactions or interactions; and $\Phi(t)$ is energy lost through heat, rupture, or stress.
A bilayer that bends without breaking, or slows the flow of ions without fully blocking them, may minimize $\Phi(t)$ and help keep $y^{\prime}(t)$ positive, even in a volatile environment.
To test this, different bilayers (made from lipids with varying tail lengths and flexibility) can be exposed to simulated early-Earth conditions: heating and cooling, dehydration, solute fluctuations. The prediction is that some bilayers will consistently survive more intervals due to the way they better align with the flow of energy around them.
If this is confirmed, it would mean membranes were not just chemically possible, they were thermodynamically favored. Selection happened not by purpose, but through physics.
The persistence model predicts that lipid bilayer configurations will be filtered based on their ability to manage environmental energy fluctuations:
  • Bilayers with optimal fluidity and curvature resist rupture, minimizing $\Phi(t)$.
  • Excessively rigid or leaky membranes dissipate energy quickly and fail to persist.
  • Certain lipid compositions (e.g., unsaturated tail mixtures) promote persistence across temperature intervals.
  • Thermodynamically stable bilayers will accumulate over time even in the absence of replication.
Tests
  • Context/Gap: Bilayer membranes are central to protocell models, yet most analyses assume steady conditions rather than fluctuating environments.
  • Model Action: The thermodynamic persistence model applies the augmented inequality to bilayers under hydration and thermal cycling, with curvature, fluidity, and composition altering $y(t)$ and $\Phi(t)$.
  • Prediction (Expectation): Bilayers with optimal curvature and fluidity are predicted to resist rupture and minimize $\Phi(t)$, while rigid or leaky bilayers should collapse. Membranes enriched in unsaturated fatty acids should persist more reliably across intervals, leading to the accumulation of stable bilayers.
  • Falsification: If bilayer persistence is unaffected by curvature, fluidity, or composition - such that rigid and leaky bilayers persist as well as optimally fluid ones - the framework would be falsified in this domain.
8.5 Interpretation of outcomes
Outcome pattern Meaning for the model Consequence
All tests affirmed
(observations match predictions across amphiphile, surface, folding, micelle, and bilayer systems) Confirms that the persistence inequality $y^{\prime}(t)\geq 0$ is a valid and general descriptor of prebiotic selection. The persistence inequality equation provides a unified thermodynamic basis for abiogenesis; the next step is quantitative parameterization and planetary modeling.
Some affirmed, some null Suggests the core mechanism is partly correct but system-specific. Energy-entropy coupling governs persistence only within defined regimes of diffusion and confinement. Guides refinement of the $\Phi(t)$ penalty term and identification of environmental boundary conditions for prebiotic stability.
No tests affirmed Indicates that persistence cannot be explained by energy alignment alone; the inequality fails as a predictive criterion. The persistence model is falsified as a general theory. Either alternative variables dominate, or selection requires kinetic heredity from the outset.
8.6 Scientific risk
By stating in advance what would disprove the model, the persistence model satisfies the criterion of scientific falsifiability. The framework does not seek confirmation but exposure: if nature demonstrates that none of the defined persistence signatures occur, the theory must be abandoned or fundamentally revised. Conversely, consistent affirmation across diverse systems would strengthen the case that thermodynamic selection predates biological evolution.